\providecommand{\bysame}{\leavevmode\hbox to3em{\hrulefill}\thinspace}
\providecommand{\MR}{\relax\ifhmode\unskip\space\fi MR }
% \MRhref is called by the amsart/book/proc definition of \MR.
\providecommand{\MRhref}[2]{%
  \href{http://www.ams.org/mathscinet-getitem?mr=#1}{#2}
}
\providecommand{\href}[2]{#2}
\begin{thebibliography}{10}

\bibitem{anupindiLinearComplexitySequences2021}
Vishnupriya Anupindi and L\'{a}szl\'{o} M\'{e}rai, \emph{Linear complexity of
  some sequences derived from hyperelliptic curves of genus 2}, Cryptogr.
  Commun. \textbf{14} (2022), no.~1, 117--134.

\bibitem{cohen2005handbook}
Henri Cohen, Gerhard Frey, Roberto Avanzi, Christophe Doche, Tanja Lange, Kim
  Nguyen, and Frederik Vercauteren (eds.), \emph{Handbook of elliptic and
  hyperelliptic curve cryptography}, Discrete Mathematics and its Applications
  (Boca Raton), Chapman \& Hall/CRC, Boca Raton, FL, 2006.

\bibitem{Galbraith-book}
Steven~D. Galbraith, \emph{Mathematics of public key cryptography}, Cambridge
  University Press, Cambridge, 2012.

\bibitem{grant1990formal}
David Grant, \emph{Formal groups in genus two}, J. Reine Angew. Math.
  \textbf{411} (1990), 96--121.

\bibitem{Lange_Stein_MR1895585}
Christian G\"{u}nther, Tanja Lange, and Andreas Stein, \emph{Speeding up the
  arithmetic on {K}oblitz curves of genus two}, Selected areas in cryptography
  ({W}aterloo, {ON}, 2000), Lecture Notes in Comput. Sci., vol. 2012, Springer,
  Berlin, 2001, pp.~106--117.

\bibitem{hartshorneAlgebraicGeometry1977}
Robin Hartshorne, \emph{Algebraic geometry}, Springer-Verlag, New
  York-Heidelberg, 1977, Graduate Texts in Mathematics, No. 52.

\bibitem{hindryDiophantineGeometryIntroduction2000}
Marc Hindry and Joseph~H. Silverman, \emph{Diophantine geometry}, Graduate
  Texts in Mathematics, vol. 201, Springer-Verlag, New York, 2000, An
  introduction.

\bibitem{Kob_CM_MR1243654}
Neal Koblitz, \emph{C{M}-curves with good cryptographic properties}, Advances
  in cryptology---{CRYPTO} '91 ({S}anta {B}arbara, {CA}, 1991), Lecture Notes
  in Comput. Sci., vol. 576, Springer, Berlin, 1992, pp.~279--287.

\bibitem{Koblitz-book}
\bysame, \emph{Algebraic aspects of cryptography}, Algorithms and Computation
  in Mathematics, vol.~3, Springer-Verlag, Berlin, 1998, With an appendix by
  Alfred J. Menezes, Yi-Hong Wu and Robert J. Zuccherato.

\bibitem{lachaud2015}
Gilles Lachaud and Robert Rolland, \emph{On the number of points of algebraic
  sets over finite fields}, J. Pure Appl. Algebra \textbf{219} (2015), no.~11,
  5117--5136.

\bibitem{lange2005koblitz}
Tanja Lange, \emph{Koblitz curve cryptosystems}, Finite Fields Appl.
  \textbf{11} (2005), no.~2, 200--229.

\bibitem{ShparlinskiLange05}
Tanja Lange and Igor Shparlinski, \emph{Collisions in fast generation of ideal
  classes and points on hyperelliptic and elliptic curves}, Appl. Algebra
  Engrg. Comm. Comput. \textbf{15} (2005), no.~5, 329--337.

\bibitem{langeCertainExponentialSums2005}
Tanja Lange and Igor~E. Shparlinski, \emph{Certain exponential sums and random
  walks on elliptic curves}, Canad. J. Math. \textbf{57} (2005), no.~2,
  338--350.

\bibitem{lange2007distribution}
\bysame, \emph{Distribution of some sequences of points on elliptic curves}, J.
  Math. Cryptol. \textbf{1} (2007), no.~1, 1--11.

\bibitem{MeidlWinterhof}
Wilfried Meidl and Arne Winterhof, \emph{Linear complexity of sequences and
  multisequences}, Handbook of Finite Fields, CRC Press, Boca Raton, 2013.

\bibitem{LM_EC_FEG_MR3649088}
L\'{a}szl\'{o} M\'{e}rai, \emph{On pseudorandom properties of certain sequences
  of points on elliptic curve}, Arithmetic of finite fields, Lecture Notes in
  Comput. Sci., vol. 10064, Springer, Cham, 2016, pp.~54--63.

\bibitem{Muller_MR1648949}
Volker M\"{u}ller, \emph{Fast multiplication on elliptic curves over small
  fields of characteristic two}, J. Cryptology \textbf{11} (1998), no.~4,
  219--234.

\bibitem{mumfordTataLecturesTheta2007a}
David Mumford, \emph{Tata lectures on theta. {II}}, Modern Birkh\"{a}user
  Classics, Birkh\"{a}user Boston, Inc., Boston, MA, 2007, Jacobian theta
  functions and differential equations, With the collaboration of C. Musili, M.
  Nori, E. Previato, M. Stillman and H. Umemura, Reprint of the 1984 original.

\bibitem{Nied1}
Harald Niederreiter, \emph{Linear complexity and related complexity measures
  for sequences}, Progress in cryptology---{INDOCRYPT} 2003, Lecture Notes in
  Comput. Sci., vol. 2904, Springer, Berlin, 2003, pp.~1--17.

\bibitem{Smart_MR1685175}
N.~P. Smart, \emph{Elliptic curve cryptosystems over small fields of odd
  characteristic}, J. Cryptology \textbf{12} (1999), no.~2, 141--151.

\bibitem{Solinas_MR1759617}
Jerome~A. Solinas, \emph{Efficient arithmetic on {K}oblitz curves}, vol.~19,
  2000, Towards a quarter-century of public key cryptography, pp.~195--249.

\bibitem{stichtenothAlgebraicFunctionFields2009}
Henning Stichtenoth, \emph{Algebraic function fields and codes}, second ed.,
  Graduate Texts in Mathematics, vol. 254, Springer-Verlag, Berlin, 2009.

\bibitem{Winterhof2010}
Arne Winterhof, \emph{Linear complexity and related complexity measures},
  Selected topics in information and coding theory, Ser. Coding Theory
  Cryptol., vol.~7, World Sci. Publ., Hackensack, NJ, 2010, pp.~3--40.

\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{amsmath,amssymb,amsbsy,amsfonts,amsthm,latexsym,mathabx,
            amsopn,amstext,amsxtra,euscript,amscd,stmaryrd,mathrsfs,
            cite,array,mathtools,enumerate}

\usepackage{url}
\usepackage[colorlinks,linkcolor=blue,anchorcolor=blue,citecolor=blue,backref=page]{hyperref}

\usepackage{color}
\usepackage{a4wide}

\usepackage{color}
\usepackage{float}

\hypersetup{breaklinks=true}

\usepackage[english]{babel}

\usepackage{mathtools}
\usepackage{todonotes}

\usepackage[norefs,nocites]{refcheck}

\usepackage{enumitem}

\allowdisplaybreaks

\usepackage{todonotes}
\usepackage{url}
\usepackage[colorlinks,linkcolor=blue,anchorcolor=blue,citecolor=blue,backref=page]{hyperref}
\usepackage[noabbrev,capitalize]{cleveref}

\usepackage{tikz}
\usetikzlibrary{arrows.meta}
\usepackage{tikz-cd}

\newtheorem{theorem}{Theorem}
\newtheorem{lemma}[theorem]{Lemma}

\newtheorem{prop}[theorem]{Proposition}

\numberwithin{equation}{section}
\numberwithin{theorem}{section}

\newcommand{\Z}{\mathbb{Z}}

\newcommand{\U}{U}

\newcommand{\X}{Z}

\newcommand{\Fq}{\mathbb{F}_q}
\newcommand{\Fqn}{\mathbb{F}_{q^n}}
\newcommand{\Fqp}{\mathbb{F}_{q^n}}
\newcommand{\Fqm}{\mathbb{F}_{q^m}}
\newcommand{\Fqb}{\overline{\mathbb{F}}_q}

\newcommand{\Bf}[1]{\boldsymbol{#1}}
\newcommand{\mb}{\Bf {m}}
\newcommand{\f}{f}
\newcommand{\F}{F}
\newcommand{\qf}{\mathfrak{q}}

\def\cO{{\mathcal O}}

\def\cR{{\mathcal R}}
\def\cS{{\mathcal S}}

\title[HEC Koblitz]{Linear complexity of sequences on Koblitz curves of genus 2}

\author[V.~Anupindi]{Vishnupriya Anupindi}
\address{Johann Radon Institute for Computational and Applied Mathematics, Austrian Academy of Sciences,  Altenberger Stra\ss e 69, A-4040 Linz, Austria} 
\email{vishnupriya.anupindi@oeaw.ac.at}

\begin{document}
\keywords{elliptic curve, hyperelliptic curve, linear complexity, frobenius generator}

\subjclass[2020]{11G05, 11G20, 11K45, 11T71 }
\maketitle

\begin{abstract}
	In this paper, we consider the hyperelliptic analogue of the Frobenius endomorphism generator and show that it produces sequences with large linear complexity on the Jacobian of genus $ 2 $ curves.
\end{abstract}

\section{Introduction}
An important operation in elliptic curve based cryptosystems is to compute scalar multiples of a given group element. The standard method for computing scalar multiples is the \textit{double-and-add-method}, but faster methods have been suggested by using the Frobenius endomorphism on special curves known as Koblitz curves, see \cite{Kob_CM_MR1243654, Muller_MR1648949, Smart_MR1685175, Solinas_MR1759617}. The ideas for fast computation of scalar multiples on elliptic Koblitz curves have been generalized to hyperelliptic curves of genus $2$, see \cite{Lange_Stein_MR1895585}. 

In \cite{ShparlinskiLange05}, Lange and Shparlinski investigated the problem of choosing random elements from elliptic and hyperelliptic curves, see also \cite{lange2007distribution, LM_EC_FEG_MR3649088}. One can choose such elements by computing random scalar multiples of an initial element fixed in advance. However, Lange and Shparlinski \cite{ShparlinskiLange05}, by taking advantage of fast computation of scalar multiplication on Koblitz curves, introduced a more efficient and direct way to obtain random-looking elements, called \emph{Frobenius endomorphism generator}. 

In this paper, we study some properties of pseudorandomness of sequences derived from hyperelliptic curves of genus $ 2 $ using the Frobenius endomorphism generator. In particular, we investigate the level of randomness of such sequences in terms of linear complexity.
We recall, that the \emph{linear complexity} of a sequence $(s_n)$ of length $N$ over the finite field $\Fq$ is defined as the smallest non-negative integer $L$ such that the first $ N $ terms of the sequence $ (s_n) $ can be generated by a linear recurrence relation over $ \Fq $ of order $ L $, i.e.\ there exist $ c_0, c_1, \dots , c_{L-1} \in \Fq $ such that 
\begin{equation*}
s_{n+L} = c_0s_{n} + c_1s_{n+1} + \dots  + c_{L-1}s_{n+L-1},\quad  0 \leq n \leq N-L-1.
\end{equation*}
The linear complexity measures the unpredictability of a sequence, hence for applications in cryptography, a large linear complexity is desired. However, a large linear complexity is not a sufficient condition for the unpredictability of a sequence. For more details, see \cite{MeidlWinterhof,Nied1, Winterhof2010}. 

In \cref{sec:HEC}, we recall some properties of hyperelliptic curves and in \cref{sec:Kob_Curve}, we define the Frobenius endomorphism generator and state the main result. In \cref{sec:prep}, we collect auxiliary results which are used in the proof. In particular, we recall the Grant representation \cite{grant1990formal} of the Jacobian of a hyperelliptic curve of genus $ 2 $ and some results from \cite{anupindiLinearComplexitySequences2021}. Finally, in \cref{sec:proof}, we prove the main result. 

\section{Hyperelliptic curves}\label{sec:HEC}
Let $ \Fq $ be a finite field with characteristic $ p \geq 3 $ and $ \Fqp $ be an extension field of $ \Fq $ with $ n \geq 1 $. Let $\Fqb$ be the algebraic closure of $\Fq$. 
\subsection{Points on hyperelliptic curves}
Let $ C $ be a hyperelliptic curve of genus $ g \geq 1 $ defined over the base field $ \Fq $ by
\begin{equation}\label{def:hec}
C: Y^2 = h(X),
\end{equation}
where $ h(X) \in \Fq[X]$ is a polynomial of degree $ 2g+1 $.  For details on hyperelliptic curves, see \cite{cohen2005handbook,Galbraith-book, Koblitz-book}.
We denote the $\Fqp$-rational points of $C$ by $C(\Fqp)$, which are the solutions over $\Fqp$ of the defining equation \eqref{def:hec} together with a point $\cO$ at infinity. 
By the Hasse-Weil bound \cite[Theorem~5.2.3]{stichtenothAlgebraicFunctionFields2009}, we have
\begin{equation}\label{eq:number_of_points}
\, |\, |C(\Fqp)| -(q^n+1) |\leq 2g q^{n/2}.
\end{equation} 
%The additive group associated to curve $ C $ that we use in cryptography is the \emph{Jacobian} $J_C$ of the curve $C$, which is an abelian variety of dimension $ g $ \cite[Theorem~A.8.1.1]{hindryDiophantineGeometryIntroduction2000}. For elliptic curves (curves with genus $ g=1 $), the Jacobian $ J_C $ is isomorphic to the curve $ C $, but this is not the case for genus $ g \geq 2 $.  For higher genus hyperelliptic curves $ (g \geq 2) $, we can describe elements of the Jacobian $ J_C $ in the following way. 
\subsection{Jacobian of hyperelliptic curves}
For an affine point $P=(x,y)\in C$, we write $-P=(x,-y)$ and $-\cO=\cO$ for the point at infinity. A \emph{divisor} $D$ of $C$ is an element of the free abelian group over the points of $C$, e.\ g.\ $D =\sum_{P\in C } n_P P$  with $n_P\in \mathbb{Z}$ and $n_P = 0$ for almost all points $P$. A \emph{reduced divisor} is given by 
\begin{equation}\label{eq:reduced_div}
D = P_1+\dots +P_r-r\cO,
\end{equation}
where $1\leq r \leq g$, $P_1, \dots, P_r\in C$, $P_i\neq \cO $ for $ 1\leq i \leq r$ and $P_i\neq -P_j$ for $1\leq i<j\leq r$. 

The \emph{Jacobian} $J_C$ of the curve $C$ is the set of reduced divisors. One can define an addition operation on the set of reduced divisors, denoted by $+$, with the identity element $\cO$, which makes $J_C$ into a group.
The elements of the curve $C(\Fq)$ are represented in the Jacobian by the set
\begin{equation}\label{eq:Theta}
\Theta(\Fq)=\{D\in J_C(\Fq): D=P-\cO, P\in C(\Fq) \} \cup \{\cO\} .
\end{equation}
We also write $\Theta=\Theta(\Fqb)$.

The Frobenius endomorphism $ \sigma: \Fqb \rightarrow \Fqb, x \mapsto x^q $, extends naturally to points on $ C $, where $ \sigma((x,y)) = (x^q,y^q)  $ and $ \sigma(\cO) = \cO $. For $ D = \sum_{i=1}^{r} P_i - r \cO  \in J_C $, define $ \sigma(D) = \sum_{i=1}^{r} \sigma(P_i) - r \cO $. An element $D \in J_C$ as given in \eqref{eq:reduced_div} is said to be defined over $\Fq$ if $ \sigma(D) $ permutes the set $\{P_1,\dots, P_r\}$. We use $J_C(\Fq)$ to denote the set of elements of $J_C$ which are defined over $\Fq$. 

The characteristic polynomial, $ \chi_C (T) $ of the Frobenius endomorphism $ \sigma $ is a degree $ 2g $ polynomial with integer coefficients of the following form
\begin{equation}\label{eq:char_poly_Frob}
\chi_C (T) = T^{2g} + s_1T^{2g-1} + \dots + s_gT^g + \dots + s_1q^{g-1}T + q^g, s_i \in \Z.
\end{equation}
It follows from the Hasse-Weil Theorem \cite[Theorem 5.1.15 and 5.2.1]{stichtenothAlgebraicFunctionFields2009}, that the complex roots $ \tau_i $ of $ \chi_C $ have absolute value $ |\tau_i| = q^{1/2}, i = 1, \dots, 2g $.  For any extension degree $ n $, the cardinality of $ J_C(\Fqn) $  is given by 
\begin{equation}\label{eq:card_Jac_roots}
|J_{C}(\Fqn)| = \prod_{i=1}^{2g} (1-\tau_i ^n).
\end{equation}
In particular, we have
\begin{equation}\label{eq:size_of_jac}
(q^{n/2}-1)^{2g}\leq |J_{C}(\Fqn)|\leq (q^{n/2}+1)^{2g}, ~  n\geq 1.
\end{equation}
\subsection{Mumford representation}
A compact representation of elements of the Jacobian $ J_C $ is given by the \emph{Mumford representation} \cite{mumfordTataLecturesTheta2007a} using a pair of polynomials $ [u,v] \in \Fq[X] \times \Fq[X]$. For a reduced divisor $ D = \sum_{i=1}^{r} P_i - r \cO $ with $ P_i = (x_i,y_i) $ the Mumford representation is given by $ u =\prod_{i=1}^r(X-x_i)$ and $v$ such that $v$ interpolates the points $ P_i $ respecting multiplicities. In particular,
\renewcommand{\theenumi}{\alph{enumi}}
\begin{enumerate}
	\item \label{mum-a} $u$ is monic,
	\item $u$ divides $f-v^2$,
	\item $\deg(v)< \deg(u)\leq g$.
\end{enumerate}
%$ u $ is monic, $ v $ is either $ 0 $ or $ \deg(v) < \deg(u) \leq g$ and $ u $ divides $ f - v^2 $. 
For genus $ g=2 $, a generic element $ D = P_1 + P_2 - 2\cO $ is represented by the polynomials 
\begin{equation}\label{eq:Mumford_rep}
u = x^2 + u_1x + u_0,~ v= v_1x + v_0,~ u_i, v_i \in \Fqn, i \in \{0,1\} \text{ such that } D=[u,v].
\end{equation} 

\section{Koblitz curves and fast generation of elements in Jacobian} \label{sec:Kob_Curve}
By a hyperelliptic \emph{Koblitz curve}, we refer to a hyperelliptic curve that is defined over a small finite field and is considered over a large extension field. In this work, we avoid fields with characteristic $ 2 $ for technical reasons. For Koblitz curves, it is recommended to choose base fields $ q \leq 7 $ for computational advantage, see \cite{lange2005koblitz}. However, we do not impose this restriction for our result. 
%In this work, we choose $ 3 \leq q \leq 7 $. Typically, we choose $ \Fq, \old{q \leq 7} $ and $ \Fqn $ with $ n $ large, \warn{i.e. $ n \sim 10^m, m\geq 2 $.} For more details, see \cite{lange2005koblitz}. 

For fast generation of elements in the Jacobian $ J_C(\Fqn) $, Lange and Shparlinski in \cite{ShparlinskiLange05} introduced the following method using the Frobenius endomorphism. Here we restrict ourselves to the genus $ 2 $ case.
Let 
\begin{equation*}
\cR = \{ 0, \pm 1, \dots, \pm (q^2 -1)/2 \}
\end{equation*}
represent the set $\Z /q^2\Z$. Let $D \in J_C(\Fqn)$ be an element of order $\ell$. For fixed $k \leq n$, consider the element of $J_C$ defined as follows:
\begin{equation}\label{eq:Dm_def}
D_{\mb} = \sum_{j=0}^{k-1} m_j \sigma^{j}(D), \quad \mb = (m_0, \dots,m_{k-1}) \in \cR^k. 
\end{equation}
%In the above construction, collisions can occur, that is, different multi-indices can yield the same element. However in \cite{ShparlinskiLange05}, the authors started to investigate the randomness properties of $ D_{\mb} $. In particular, they showed that $ D_{\mb} $ is well distributed, namely, there are few collisions. For elliptic curves, the randomness properties of the points $  D_{\mb} $ were investigated in \cite{lange2007distribution} and \cite{LM_EC_FEG_MR3649088}. \warn{More content to be added...}
It is natural to expect that the divisors $ D_{\mb} $ defined by \eqref{eq:Dm_def} are sufficiently uniformly distributed. Lange and Shparlinski \cite{ShparlinskiLange05} showed that $ D_{\mb} $ do not take the same value too often (which would otherwise have catastrophic implications for their cryptographic applications).

In this paper we further investigate the randomness properties of $ D_{\mb} $. Namely, we show that different statistics of the divisors $ D_{\mb} $, like the Mumford coordinates $ u_i, v_i $ as in \eqref{eq:Mumford_rep}, possess large linear complexity if the divisors $ D_{\mb} $ are arranged in a natural way, say in lexicographic ordering. More precisely, let 
$\f \in \Fqp(J_C)$ be a rational function in the function field of the Jacobian.  We arrange the elements of $\cR^k$ with a lexicographic ordering and define the sequence $ (w_{\mb})_{\mb \in \cR^k} $ with
\begin{equation}\label{eq:seq_def}
w_{\mb} = \begin{cases}
f(D_{\mb}) &\text{if $D_{\mb}$ is not a pole of $f$,} \\
0  &\text{otherwise.}
\end{cases} 
\end{equation}
%We also define $ \mb + i $ as the $ i^{\text{th}} $ element after $ \mb $ with respect to the lexicographic ordering. 
Throughout the paper, $U \ll V $ is equivalent to the inequality $ |U| \leq cV $ with some constant $ c > 0 $.
Our main result is the following bound on the linear complexity of $ (w_{\mb})_{\mb \in \cR^k} $.
\begin{theorem}\label{thm:main_result}
	Let $ C $ be a hyperelliptic curve of genus $ 2 $, defined over the base field $ \Fq $ and let $J_C(\Fqn)$ be its Jacobian over the extension field $ \Fqn $. Let the characteristic polynomial of the Frobenius endomorphism $\chi_C$ be irreducible. Let $f \in \Fqn(J_C)$ be a rational function with pole divisor of the form $\alpha \Theta, \alpha \in \Z, \alpha \geq 1 $. If $D \in J_C(\Fqn)$ is of prime order $\ell, \ell \nmid q^2$, then for any $ k $ where $ 1 \leq k \leq n $ with $(w_{\mb})_{\mb \in \cR^k}$ as defined in \eqref{eq:seq_def}, we have
	\begin{equation}\label{eq:thm_main}
	L(w_{\mb}) \gg  \frac{\min\{ q^{3k/2},\ell/q^8 \}}{ q^n \deg \f }.
	\end{equation} 
\end{theorem}
The result is non-trivial if $ k \geq 2n/3 $ and $ \ell \geq q^{n+8} $. In the ideal case, $k=n$, $ \deg \f =1 $ and $ \ell \sim q^{2n} $, we obtain $ L(w_{\mb}) >  c q^{n/2} $ for some constant which may depend on $ \deg \f $. Examples for rational functions with $ \deg \f =1 $ are the Mumford coordinates \eqref{eq:Mumford_rep}.

We assume the characteristic polynomial of Frobenius endomorphism $ \chi_C$ to be irreducible, in particular, $ \chi_C $ is irreducible over $ \Z $. Practically, this is the most interesting case, since, by \eqref{eq:card_Jac_roots} any non-trivial factor of $ \chi_C $ leads to a non-trivial factor of the group order, which we want to avoid. 

We remark, that in \eqref{eq:Dm_def}, if we replace the Frobenius map $ \sigma $ with the multiplication map $ [2]: D \mapsto 2D $, and if we use colexicographic ordering for arranging sequence elements $ D_{\mb} $, then we are in the linear congruential generator case, for which we proved a stronger bound in \cite{anupindiLinearComplexitySequences2021}.

We also remark, that Lange and Shparlinski \cite{ShparlinskiLange05, lange2007distribution} defined and investigated the randomness properties of similar, but not completely analogous point-set for the elliptic curve case. Later, M{\'e}rai \cite{LM_EC_FEG_MR3649088} studied the randomness properties of a sequence of elements from this point set,  by arranging elements in a sequence using lexicographic ordering.

The proof of \cref{thm:main_result} is based on the method of \cite{LM_EC_FEG_MR3649088}, the results of \cite{ShparlinskiLange05} and taking advantage of the explicit addition formulas for genus $ 2 $ provided by Grant \cite{grant1990formal}. 

%This is indeed the case when $ \f $ is one of the Mumford coordinates, where $ \deg \f =1$.
% for $ \f = u_0 = -z_{12}, u_1 = -z_{22}, v_0 = z_{122}, v_1 = z_{222} $. \commV{Not perfect %yet, $z_{ijk}$ not defined at this point.}

\section{Preparation}\label{sec:prep}
The aim of this section is to collect some technical results for the proof of the main theorem. We use the \emph{Grant representation} of a hyperelliptic curve of genus $ 2 $ since it provides explicit addition formulas. This allows us to prove the degree estimate in \cref{prop:F_nonconst}. 
%In this section, we will recall the \emph{Grant representation} \new{of a hyperelliptic curve of genus $ 2 $} and some other technical lemmas from \cite{anupindiLinearComplexitySequences2021}. The aim of this section is to prove \cref{prop:F_nonconst}, which is a key ingredient for proof of the main theorem.

\subsection{Arithmetic for genus 2 using Grant representation}
In order to implement the group law $(Q,R) \mapsto Q+R$ on the Jacobian, one can use Cantor's algorithm which uses the Mumford representation. However, this algorithm is implicit.
%Algorithms based on Cantor's algorithm use the Mumford representation to implement the group law $(Q,R) \mapsto Q+R$ on the Jacobian. 
In this work, we use the explicit addition formulas provided by Grant \cite[Theorem 3.3]{grant1990formal}.

Let $C$ be the hyperelliptic curve of genus $g = 2$ defined by \eqref{def:hec} with
$$
h(X) = X^5 + b_1X^4 + b_2X^3 + b_3X^2 + b_4X + b_5 \in \Fq[X],
$$
for the finite field $\Fq$ with characteristic $p \geq 3 $.
In \cite{grant1990formal}, Grant provides an embedding of $J_C$ into the projective space $\mathbb{P}^8$.

Let 
\begin{equation} \label{eq:poly_ring}
\Fq[\mathbf{\X}] = \Fq[\X_{11},\X_{12},\X_{22},\X_{111},\X_{112},\X_{122},\X_{222},\X]
\end{equation}
be a polynomial ring over $\Fq$ in $8$ variables. The following proposition gives us a set of defining equations for the Jacobian, see \cite[Corollary 2.15]{grant1990formal}.

\begin{prop}
	There are polynomials $ f_1, \dots , f_{13} \in \Fq[\mathbf{\X}] $ 
	such that
	\begin{equation*}
	J_C \cong V(f_1^h,\dots,f_{13}^h) = \{ z \in \mathbb{P}^8 : f_i^h(z) = 0, 1 \leq i \leq 13 \},
	\end{equation*}
	where $f_i^h$ denotes the homogenized polynomial with respect to the variable $\X_0$.
	Moreover, an embedding $\iota : J_C \rightarrow \mathbb{P}^8$ is given by
	\begin{equation}\label{eq:iota}
	\iota(D) = 
	\begin{cases}
	(1:z_{11}:z_{12}:z_{22}:z_{111}:z_{112}:z_{122}:z_{222}:z)  & \text{if~} D \in J_C\setminus \Theta, \\ 
	(0:0:0:0:1:0:0:0:0) & \text{if~} D = \cO ,\\
	(0:0:0:0:-x^3: - x^2 : -x : 1 : -y) & \text{if~} D = P - \cO \in \Theta \setminus\cO,
	
	\end{cases}
	\end{equation}
	where $ P = (x,y) $.
\end{prop}
See \cref{sec:def_eq_J} for the polynomial expressions of $f_1,\dots, f_{13}$ in the same notation as used in this work. For $D = (x_1,y_1) + (x_2,y_2) - 2\cO \in J_C(\Fq)\setminus \Theta(\Fq)$, the components $z_{jk}, z_{jkl}$ of $\iota(D)$ can be expressed as rational functions in the coordinates $(x_1,y_1)$ and $(x_2,y_2)$. 

We denote the affine part of $J_C$ with respect to variable $\X_0$ under $\iota$ by $U$. Then \\ $ U = J_C\setminus \Theta.$
Moreover, 
by \cite[Theorem~2.5]{grant1990formal}, we have
\begin{equation}\label{eq:def_U}
\U \cong V(f_1,\dots, f_6)
\end{equation}
Since $ J_C$ is irreducible and has dimension $ 2 $, it follows that $ U $
is irreducible, dense, and has dimension $ 2 $, see \cite[Example 1.1.3 ]{hartshorneAlgebraicGeometry1977}. 
As a result, $ \Fq(U) = \Fq(J_C)$, see \cite[Theorem~3.4]{hartshorneAlgebraicGeometry1977}.

For a rational function $ h \in \Fq(U) $, we define its degree by choosing a representative element $\frac{h_1}{h_2}$ of the equivalence class $ h $, such that $\deg h_1$ is minimal and set 
$$
\deg h =\max\{\deg h_1, \deg h_2\}.
$$

We summarize the algebraic properties of the group law in the Grant representation. For explicit expressions, see \cref{sec:addition_formulas}.
\begin{lemma}\label{lem:Grant_q}
	Assume that $Q,R,Q+R,Q-R \in U$. Let 
	\begin{equation}\label{eq:q}
	\qf(Q,R) = z_{11}(Q)-z_{11}(R)+z_{12}(Q)z_{22}(R)-z_{12}(R)z_{22}(Q).
	\end{equation}
	Then there are explicit formulas for $ z_{jk}(Q+R), z_{jkl}(Q+R) $ which are rational functions in $ z_{jk}(Q),z_{jk}(R), z_{jkl}(Q) z_{jkl}(R) $ and $ \qf(Q,R) $ for $ 1 \leq j \leq k \leq l \leq 2 $.
\end{lemma}

We recall \cite[Lemma 2.3]{anupindiLinearComplexitySequences2021} which will be used in \cref{prop:F_nonconst}.

\begin{lemma}\label{lem:Grant_q_dim}
	Assume that $Q,R,Q+R,Q-R \in U$. Let $ \qf(Q,R) $ be defined by \eqref{eq:q} and set $\qf_R(Q)=\qf(Q,R) $.
	Then for any fixed $R\in \U$,  the zero set $\{ \qf_R(Q)=\qf(Q,R)=0 \}$ has dimension one and $\Theta \pm R\subset\{\qf_R=0\}$. Moreover
	if $R' \in U$ with $R\neq \pm R'$, then 
	\begin{equation}\label{eq:cardinality_zeros_q_1}
	|\{\qf_{R} = 0\} \cap \{\qf_{R'} = 0\} \cap \U| \leq 20.
	\end{equation}
\end{lemma}

One can show that for $D \in U(\Fqm)$, where $ m \in \Z, m \geq1 $, we have $ |\{\Theta(\Fqm) + D\} \cap \Theta(\Fqm)| \leq 2 $. See \cite[Lemma~2.4]{anupindiLinearComplexitySequences2021}. Thus
\begin{equation}\label{eq:Theta_intersection}
|\{\Theta(\Fqm) + D\} \, \cap \,U| \geq |\Theta(\Fqm)|-2.
\end{equation}

\begin{prop}\label{prop:F_nonconst}
	Let $\f \in \Fqp(U)$ be a rational function with a pole divisor of the form $\alpha \Theta, \alpha \in \Z, \alpha \geq 1 $. Let $L$ be a positive integer and 
	%\begin{equation}\label{eq:bound_L}
	%L < \frac{|\Theta(\Fqp)| - 2}{20}. 
	%\end{equation}
	let $R_0, \dots, R_L \in J_C(\Fqp)$ such that $R_i \notin \Theta(\Fqp)$ and $R_L \neq \pm R_j$ for $ 0 \leq i \leq L $ and $ 0 \leq j \leq L-1$. Let $c_0. \dots, c_L \in \Fqp$ with $c_L \neq 0$. Then the rational function $\F \in \Fqp(U)$, with 
	\begin{equation*}
	\F(Q) = \sum_{i=0}^{L} c_i \f(Q + R_i )
	\end{equation*}
	is non-constant and has degree 
	\begin{equation}\label{eq:deg_F}
	\deg \F \leq 6(L+1) \deg \f.
	\end{equation}
\end{prop}
\begin{proof}
	Defining the function $\f_{R_i} : Q \mapsto \f(Q + R_i )$ yields 
	$$
	\F(Q) = \sum_{i=0}^{L-1} c_i f_{R_i}(Q) + c_L f_{R_L}(Q).
	$$ 
	To prove that $ \F $ is non-constant, we show that there exists $ Q \in U $ such that it is a pole of $ \f_{R_L} $, but not a pole of any other terms $ \f_{R_i}$ for $ i < L $.
	
	Observe that $\f_{R_L} $ has a pole at $ Q $ when $ Q \in \Theta - R_L $, in particular, when $ Q \in \Theta(\Fqm) - R_L $, for $ m \geq 1 $ independent of $ n $. Define $\qf_{R_i} = \qf(Q,R_i)$. 
	From \cref{lem:Grant_q_dim}, we know that $ \Theta(\Fqm) -R_i \subseteq \{\qf_{R_i} = 0\} $.  Hence, by \eqref{eq:cardinality_zeros_q_1} we obtain
	$$
	\left| \Big((\Theta(\Fqm) - R_L) \cap \U \Big) \cap \{\qf_{R_i} = 0\} \right|\leq
	\left| \{\qf_{R_L}=0\} \cap \{\qf_{R_i} = 0\} \cap \U\right|\leq 20 .
	$$
	Thus, by \eqref{eq:Theta_intersection}, we obtain
	\begin{align}\label{eq:set}
	&\left|\Big((\Theta(\Fqm) - R_L)\cap \U \Big) \setminus \left( \bigcup\limits_{i=0}^{L-1} \{\qf_{R_i} = 0 \} \right)\right| \\ \notag
	=&\left| \bigcap\limits_{i=0}^{L-1} \bigg( \Big((\Theta(\Fqm) - R_L) \cap \U \Big)  \setminus \{\qf_{R_i} = 0\} \bigg)\right| \geq |\Theta(\Fqm)|-2-20L.
	\end{align}
	We pick $ m $ such that $ |\Theta(\Fqm)|-2-20L~ > 0 $. Hence, there exists a point $Q$ which is a pole of $ \f_{R_L}$ but not a pole of any other term of $\F$. Hence, $\F$ is non-constant.
	
	To estimate the degree of $\F$, we first estimate the degree of $\f_{R_i}$. For arbitrary $i$, define $R = R_i$. 
	We define 
	\begin{equation*}
	z_{jk}^R(Q) = z_{jk}(Q+R),
	\quad
	z_{jkl}^R(Q) = z_{jkl}(Q+R),
	\quad
	z^R = z(Q+R).
	\end{equation*}
	Then we can write $ \f_{R}(Q) $ as 
	\begin{equation*}
	\f_{R}(Q) =\f \big(Q + R\big) =\f \big(z_{11}^{R}(Q), \dots , z_{222}^{R}(Q),z^{R}(Q)\big).
	\end{equation*}
	We can consider $z_{jk}^R(Q), z_{jkl}^R(Q)$ to be rational functions in the variables $z_{jk}(Q)$,  $ z_{jkl}(Q)$ and $z(Q)$, see \cref{sec:addition_formulas}. Then it follows from the explicit formulas of these functions that 
	\begin{equation*}
	\deg z_{jk}^R \leq 3 \text{~  ,  ~} \deg z_{jkl}^R \leq 4 \text{~  and  ~} \deg z^R \leq 6.
	\end{equation*}
	Hence we obtain 
	\begin{align*}
	\deg \f_{R} \leq (\deg \f) \big(\max \{ \deg z_{jk}^{R}, \deg z_{jkl}^{R}, \deg z^R \}\big) = 6 \deg \f,
	\end{align*}
	and thus
	\begin{equation*}
	\deg \F \leq \deg \bigg( \sum_{i=0}^{L} c_i \f_{R} \bigg) \leq 6(L+1) (\deg \f).
	\end{equation*}
	
\end{proof}

\subsection{Bounds on the number of zeros of a system of polynomial equations over a finite field.}
%While \cref{prop:F_nonconst} gives an upper bound on the degree of $\F$, this does not immediately give an upper bound on the number of zeros of $\F$ over finite field $\Fqp$, since the zero set of $\F$ has dimension $1$. Now, we give an estimate for the number of zeros of $\F$ over finite field $\Fqp$.

Let $f_1,\dots, f_k\in \Fqp[X_1,\dots, X_m]$. We denote the vanishing set of $ f_1, \dots, f_k $ over $ \Fqp $ by 
$$
V_{\Fqp}(f_1,\dots, f_k)=\{\mathbf{x}\in\Fqp^m: f_1(\mathbf{x})=\dots=f_k(\mathbf{x})=0 \}, 
$$
and the vanishing set over the algebraic closure $ \Fqb $ by 
$$
V(f_1,\dots, f_k)=V_{\Fqb}(f_1,\dots, f_k).
$$
For each $ m \in \Z, m \geq 1 $, we define affine $ m $-space over $ \Fqb $ to be 
\begin{equation}
\mathbb{A}^m (\Fqb) =  \{ (x_1, \dots, x_m) : x_i  \in \Fqb, 1 \leq i \leq m \}. 
\end{equation}
The following result gives us bounds for the cardinality of algebraic sets over finite fields,
\cite[Corollary 2.2]{lachaud2015}. 
\begin{lemma}\label{lemma:bezout}
	Let  $f_1,\dots, f_k\in \Fqp[X_1,\dots, X_m]$ such that $  V(f_1,\dots, f_k) $ has dimension $ d $ in $ \mathbb{A}^m(\Fqb) $. Then
	$$
	|V_{\Fqp}(f_1,\dots, f_k)|=| V(f_1,\dots, f_k) \cap \Fqp^{m}|\leq (q^n)^d \prod_{i=1}^k \deg f_i.
	$$
\end{lemma} 

%The following lemma gives an estimate of the number of zeros of a rational function $ \F $ over a finite field $ \Fqp, n\geq 1 $, where $ \dim V(\F) = 1 $. It is a part of the proof of \cite[Theorem~4.1]{anupindiLinearComplexitySequences2021}, but for the sake of completeness, we state the result and include the proof.

\begin{lemma}\label{lem:Zeros_F_deg}
	Let $f_1,\dots,f_6$ be the defining equations of $U$ as in \eqref{eq:def_U}, let $\F \in \Fqp(U)$ be a non-constant rational function and let $G_1 / G_2$ be a representation of $\F \in \Fqp(\mathbf{\X})$ as a rational function. Then 
	\begin{equation}\label{eq:vanishing_F}
	|V_{\Fqp}(f_1,\dots,f_6,G_1)| \leq  216 q^n \deg \F. 
	\end{equation}
\end{lemma}
\begin{proof} Since $ U $ has dimension $ 2 $ and $ \F $ is non-constant on $ U $, $ V(f_1,\dots,\f_6, G_1) $ has dimension $ 1 $ in $ \mathbb{A}^8 $. Applying \cref{lemma:bezout}, we obtain
	\begin{equation*}
	|V_{\Fqp}(f_1,\dots,f_6,G_1)| \leq q^n \deg G_1 \prod_{i=1}^6 \deg f_i  \leq 216 q^n \deg \F.
	\end{equation*} 
\end{proof}

\subsection{Linear complexity}
We recall the following result on the linear complexity, see \cite[Lemma~ 6]{langeCertainExponentialSums2005}.
\begin{lemma}\label{lemma:linComp}
	Let $(s_n)$ be a linear recurrent sequence of order $L$ over $\Fq$ defined by a linear recursion
	$$
	s_{n+L}=c_0s_n +\dots + c_{L-1}s_{n+L-1}, \quad n\geq 0.
	$$
	Then for any $T\geq L+1$ and pairwise distinct positive integers $j_1,\dots, j_{T}$, there exist $\lambda_1,\dots, \lambda_T \in \Fq$, not all equal to zero, such that
	$$
	\sum_{i=1}^T\lambda_is_{n+j_i}=0, \quad n\geq 0.
	$$
\end{lemma} 

\subsection{Number of torsion elements}
We need the following result on the number of torsion elements in the Jacobian of hyperelliptic curves over finite fields. 
% Let $ J_C $ be the Jacobian of a hyperelliptic curve of genus $ g $. Let $ m \in \Z $. An element $ D $ is an $ m-$torsion element whenever $ mD = \cO $. 
\begin{lemma}\label{lem:torsion}
	%Let $ m $ be an integer such that characteristic of the field $ \Fq $ does not divide $ m $. 
	Let $ m $ be an integer coprime to the characteristic of $ \Fq $. Then, 
	\begin{equation*}
	\# \{D \in J_C(\Fqb) : mD = \cO \} = m^{2g}.
	\end{equation*}
\end{lemma}
For a proof, we refer to  \cite[Theorem~A.7.2.7]{hindryDiophantineGeometryIntroduction2000}.

\subsection{Collisions}

We now turn our attention to the collisions which can occur in \eqref{eq:Dm_def}. Let $T_k(Q)$ be the number of $k$-tuples $\mb = (m_0,\dots,m_{k-1}) \in \cR^k $ such that $D_{\mb} = Q$.
We recall the following result from \cite[Theorem 2]{ShparlinskiLange05}, which gives an upper bound for $ T_k(Q) $. This upper bound implies that the elements generated by \eqref{eq:Dm_def} do not take the same value too often and are sufficiently uniformly distributed.

\begin{prop}\label{lem:distinct_points}
	Let $C$ be a hyperelliptic curve of genus $2$ defined over $\Fq$ such that the characteristic polynomial of the Frobenius endomorphism $\chi_C$ is irreducible. Let $D \in J_C(\Fqn)$ of prime order $\ell$. Then for any integers $k$ and $e$ with $1\leq e \leq k$ and $q^{2e} \leq (q^{1/2} -1)^4 q^{-8} \ell $, and for every element $Q \in J_C(\Fqn)$, the bound $T_k(Q)\leq q^{2(k-e)}$ holds.
\end{prop}

The bound of \cref{lem:distinct_points} shows that if $ k $ is small and $q^{2k} \leq (q^{1/2} -1)^4 q^{-8} \ell $, then all the elements $D_{\mb}$ are distinct. We observe that if $ q^{2e} \ll \ell /q^8 $ then $q^{2e} \ll (q^{1/2} -1)^4 q^{-8}\ell $. For larger $k$, choosing $e$ maximal such that $ q^{2e} \ll \ell / q^8 $ yields 
\begin{equation}\label{eq:collision_bound}
T_k(Q) \leq \max \{1,  q^{2k - 2e} \} \ll \max \Big\{ 1, \frac{q^{2k + 8}}{\ell} \Big\} .
\end{equation}

\section{Proof of the main theorem}\label{sec:proof}
Let $ \chi_C(T) = T^4 + s_1T^3 + s_2 T^2 + s_1qT + q^2 $ be the characteristic polynomial of the Frobenius endomorphism for genus $ 2 $. The following result is a crucial step in the proof of the main theorem.  
\begin{lemma}\label{lem:sigma_distinct} 
	If $D \in J_C(\Fqn) $ has prime order $\ell$, and $\ell $ does not divide the constant term of $ \chi_C $, then $\sigma(D) \neq \cO$.
	%such that $\ell$ does not divide the constant term of $\chi_C$, i.e. $\ell \nmid q^2$, then $\sigma(D) \neq \cO$.
\end{lemma}
\begin{proof}
	If $ \sigma(D) = \cO $, then by definition of $ \chi_C $, we have that
	\begin{equation*}
	q^2D = -\sigma(D)^4 - s_1\sigma(D)^3 - s_2 \sigma(D)^2 - s_1q\sigma(D)  = \cO.
	\end{equation*}
	Thus, the order $ \ell $ of $ D $ divides $ q^2 $, which is the constant term of $ \chi_C $.
\end{proof} 
%Therefore, if $D_{\mb} \neq D_{\Bf{n}}$, then $\sigma^j(D_{\mb}) \neq \sigma^j(D_{\Bf{n}}), \new{\mb,\Bf{n} \in \cR^k, \mb \neq \Bf{n} }, j \in \Z, j \geq 1$.

\begin{proof}[Proof (\cref{thm:main_result})]
	We fix $ r = \max \{  \left\lfloor \frac{k}{4} \right\rfloor , 1 \} $.
	Let $\mb \in \cR^k$, we can write 
	\begin{equation*}
	\mb = (\Bf{\mu},\Bf{\nu}), \Bf{\mu} \in \cR^{r}, \Bf{\nu}\in \cR^{k-r} .
	\end{equation*}
	Let $N_r$ and $N_{k-r}$ be the number of distinct elements $D_{\Bf{\nu}}, \Bf{\nu} \in \cR^r$ and $\Bf{\nu} \in \cR^{k-r}$ respectively. We can assume that $ \ell \gg q^{\frac{3}{4}n +8} $, since otherwise \eqref{eq:thm_main} holds trivially. Therefore, $ \max \{q^{2r}, q^{2k-2r} \} \ll \ell/q^8  $. Hence, by \eqref{eq:collision_bound}, we obtain
	%\begin{equation*}
	%N_r \geq \frac{\# \cR^{r}}{\max_Q T_r(Q)}  \gg \frac{q^{2r}}{\max \{1, q^{2r+8}/ \ell \} } = \min \Big\{q^{2r},\frac{\ell}{q^8} \Big\} 
	%\end{equation*}
	%and similarly
	\begin{equation}\label{eq:bound_Nkr}
	N_{k-r} \geq \frac{\# \cR^{k-r}}{\max_Q T_{k-r}(Q)}  \gg \frac{q^{2(k-r)}}{\max \{1, q^{2(k-r) + 8}/ \ell \} } = \min \Big\{q^{2(k-r)},\frac{\ell}{q^8} \Big\} .
	\end{equation}
	Let $L$ be the linear complexity of the sequence $(w_{\mb})_{\mb \in \cR^k}$ as defined in \eqref{eq:seq_def}. We can assume that 
	\begin{equation}\label{eq:bound_L_thm}
	L < \min\left\{N_r, \frac{|J_C(\Fqp)|-|\Theta(\Fqp)|}{|\Theta(\Fqp)|+16}\right\}, 
	\end{equation}
	since otherwise the theorem holds trivially.
	
	Since by \eqref{eq:bound_L_thm}, we assume that $ L < N_r $, there exist $ L+1 $  vectors  $\Bf{d_0},\dots,\Bf{d_L}\in \cR^r$ such that $D_{\Bf{d_0}}, \dots , D_{\Bf{d_L}}$ are distinct.
	%By \cref{lem:sigma_distinct}, we obtain that $\sigma^{k-r}(D_{\Bf{d_0}}),\dots,\sigma^{k-r}(D_{\Bf{d_L}}) $ are also distinct.
	
	We fix these vectors and for each $ j = 0, \dots,L $ define the sequence 
	\begin{equation*}
	a_{j}(\Bf{s}) = w_{(\Bf{d_j},\Bf{s})}, \quad \Bf{s} \in \cR^{k-r},
	\end{equation*} 
	%\begin{equation*}
	%\f_{j}(\Bf{s}) = f(D_{(\Bf{d_j}, \Bf{s})}), \quad \Bf{s} \in %\cR^{k-r},
	%\end{equation*}
	where again the elements $ a_{j}(\Bf{s}) $ are arranged in a sequence by using lexicographic ordering for vectors $ \Bf{s} $. The sequences $ (a_{j}(\Bf{s}) )_{\Bf{s} \in \cR^{k-r}} $ are parts of $ (w_{\mb})_{\mb \in \cR^k} $, that is, they are consecutive elements in  $ (w_{\mb})_{\mb \in \cR^k} $, as $ \Bf{s} $ runs through $ \cR^{k-r} $. By \cref{lemma:linComp}, these sequences are linearly dependent, i.e. there exist constants $ c_0, \dots . c_L \in \Fqp $, not all zero, such that 
	%Hence, they satisfy the same linear recurrence relation of order $ L $. Then these sequences are in a vector space of dimension at most $ L $ so they are linearly dependant, i.e. there exist constants $ c_0, \dots . c_L \in \Fqp $, not all zero, such that 
	\begin{equation}\label{eq:seq_LC}
	c_0  w_{(\Bf{d_0}, \Bf{s})}  + \dots + c_L  w_{(\Bf{d_L}, \Bf{s})}  = 0 ,  \quad \Bf{s} \in \cR^{k-r}.  
	\end{equation} 
	%By \eqref{eq:Dm_def}, we can rewrite this as 
	%\begin{equation*}
	%c_0  f(D_{\Bf{d_0}} + \sigma^{r}(D_{\Bf{s}}) ) + \dots + c_L  f(D_{\Bf{d_L}} + \sigma^{r}(D_{\Bf{s}}) )  = 0 ,  \quad \Bf{s} \in \cR^{k-r}. 
	%\end{equation*} 
	Note that for $ \mb = (\Bf{d_j}, \Bf{s}) $, $ D_{\mb} = D_{(\Bf{d_j}, \Bf{s})} = D_{\Bf{d_j}} + \sigma^r(D_{\Bf{s}}) $ by \eqref{eq:Dm_def}.
	
	We would like to avoid collision of elements $  D_{\Bf{d_j}} , j \in \{0,\dots, L\}$ with $  \Theta(\Fqp) $. We claim that there exists an element $ R \in J_C(\Fqn)$ such that
	\begin{align}
	D_{\Bf{d_i}}  + R \notin \Theta(\Fqp), \quad &\text{~for~} 0 \leq i \leq L, \label{eq:R_cond_1}\\
	D_{\Bf{d_L}}  + R \neq -( D_{\Bf{d_j}}  + R),\quad &\text{~for~} 0 \leq j \leq L-1. \label{eq:R_cond_2}
	\end{align} 
	We count the number of elements $ R \in J_C(\Fqn) $ such that $ R $ does not satisfy \eqref{eq:R_cond_1} or \eqref{eq:R_cond_2}. There are at most $ (L+1)|\Theta(\Fqp)| $ choices for $ R $ such that $ 		 D_{\Bf{d_i}}  + R \in \Theta(\Fqp) $ for some $ 0 \leq i \leq L $.
	%Indeed, if \eqref{eq:R_cond_1} was not satisfied, then we have at most $ (L+1)|\Theta(\Fqp)| $ choices for $ R $ such that $ 	\sigma^{k-r}(D_{\Bf{d_i}}) + R \in \Theta(\Fqp) $ for $ 0 \leq i \leq L $. 
	
	Furthermore, if \eqref{eq:R_cond_2} was not satisfied, then we obtain that 
	\begin{equation*}
	-( D_{\Bf{d_L}}  +  D_{\Bf{d_j}} ) = 2R, \text{~for some~} 0 \leq j \leq L-1. 
	\end{equation*}
	By \cref{lem:torsion}, we obtain that there are at most $ 16 $ elements $ R \in J_C(\Fqb) $ such that $ 2R = \cO $.
	%of $ n $ torsion elements of $ J_C(\Fqb) $ is $ n^{2g} $, i.e for $ g=2=n $ there are at most $ 16 $ elements $ R \in J_C(\Fqb) $ such that $ 2R = \cO $. See  \cite[Theorem A.7.2.7]{hindryDiophantineGeometryIntroduction2000}. 
	Therefore, there are at most $ 16 L $ choices for $ R $ such that $2R = -( D_{\Bf{d_L}}  +  D_{\Bf{d_j}} ) $ for some $ j \in \{0,\dots,L-1\} $.
	By \eqref{eq:bound_L_thm}, we know that $$ |J_C(\Fqp)| - (L+1)|\Theta(\Fqp)| - 16L > 0, $$ hence there exists $ R \in J_C(\Fqp) $ such that \eqref{eq:R_cond_1} and \eqref{eq:R_cond_2} are satisfied.
	
	Let $R_i =  D_{\Bf{d_i}}  + R $. 
	Consider the function
	\begin{equation}\label{eq:F_thm}
	\F(Q) = \sum_{i=0}^{L} c_i \f(Q + R_i).
	\end{equation}
	By \cref{prop:F_nonconst} we know that $\F$ is non-constant and has degree at most $  6(L+1) (\deg \f) $.
	
	%We observe that if $ f(D_{\mb}) $ is defined for $ \mb = (\Bf{d_j},\Bf{s}), \Bf{d_j} \in \cR^r,\Bf{s} \in \cR^{k-r} $, then by \eqref{eq:Dm_def}, we can write $ f(D_{\mb}) = f(D_{\Bf{d_j}} + \sigma^{r}(D_{\Bf{s}})) $.
	%Writing $ D_{\mb} = D_{\Bf{d_j}} + \sigma^{r}(D_{\Bf{s}}) $, 
	
	We observe that if $ \F $ has a pole at $ Q $, then $ Q $ must have a form $ Q = \sigma^{r}(D_{\Bf{s}})  - R, \Bf{s} \in \cR^{k-r}, $ with $ Q \in \Theta(\Fqp) $ or $ Q \in \Theta(\Fqp) \pm R_i $ for $ 0 \leq i \leq L $. Hence, defining set $ \cS $ as follows ensures that for $ Q \in \cS $, the sum in \eqref{eq:F_thm} does not contain any poles.
	Define 
	\begin{align*}
	\cS = \{ Q \in J_C(\Fqn) :~ &Q = \sigma^{r}(D_{\Bf{s}})  - R, \Bf{s} \in \cR^{k-r}, \text{ with } Q \notin \Theta(\Fqp) \\* &\text{ and }  Q \pm R_i \notin \Theta(\Fqp), \text{ for } 0 \leq i \leq L \}.
	\end{align*} 
	%The definition of set $ \cS $ ensures that for $ Q \in \cS $, the sum in \eqref{eq:F_thm} does not contain any poles. 
	Hence by \eqref{eq:seq_def} and \eqref{eq:seq_LC}, $ \F(Q) = 0 $ for $ Q \in \cS $. 
	
	Now we give a lower bound for $ |\cS| $. We observe that if $ D $ has prime order $ \ell $, then $ \sigma^j(D)$ also has order $ \ell $, since $ \sigma $ is an endomorphism and hence additive. Combining this with \eqref{eq:Dm_def}, we see that if $ \ell D_{\mb} = \cO $, then either $D_{\mb}= \cO$ or it has order $ \ell $, since $ \ell $ is prime. Hence, by \cref{lem:sigma_distinct}, we obtain that if $D_{\mb} \neq D_{\Bf{n}}$, then $\sigma^j(D_{\mb}) \neq \sigma^j(D_{\Bf{n}}), \mb,\Bf{n} \in \cR^k, j \in \Z, j \geq 1$. Therefore, the number of distinct elements $ \sigma^{r}(D_{\Bf{s}}) - R, \Bf{s} \in \cR^{k-r} $ is $ N_{k-r} $.
	We observe that for $ Q = \sigma^{r}(D_{\Bf{s}})  - R, \Bf{s} \in \cR^{k-r} $,
	\begin{equation}\label{eq:bad_Q_1}
	|\{Q \in J_C(\Fqn) :  Q \in \Theta(\Fqn) \}| \leq |\Theta(\Fqp)|
	\end{equation}
	and
	\begin{equation}\label{eq:bad_Q_2}
	|\{Q \in J_C(\Fqn) : Q \pm R_i \in \Theta(\Fqn) \}| \leq  2 (L+1)|\Theta(\Fqp)|.
	\end{equation}
	%Hence, counting the number of distinct elements $ \sigma^{r}(D_{\Bf{s}})  $ is the same as counting the number of distinct elements $  D_{\Bf{s}} $.There are at most $|\Theta(\Fqp)|$ elements $Q $ with $ Q \in \Theta(\Fqp)$ and similarly for each $i,~ 0\leq i \leq L$, there are $2|\Theta(\Fqp)|$ elements  $Q$ with $ Q \pm R_i \in \Theta(\Fqp)$. 
	Hence, by \eqref{eq:bad_Q_1} and \eqref{eq:bad_Q_2} we obtain, 
	\begin{equation}\label{eq:S_lowerbound}
	|\cS| \geq N_{k-r} - 2 (L+1)|\Theta(\Fqp)| - |\Theta(\Fqp)|.
	\end{equation}
	
	To give an upper bound for $ |\cS| $, 
	%we first observe that by \eqref{eq:bound_L_thm}, \eqref{eq:claim} is satisfied since
	%\begin{equation*}
	%|\cS| \geq N_{k-r} - 2 (L+1)|\Theta(\Fqp)| - |\Theta(\Fqp)| > 2|\Theta(\Fqp)| + 2(q^n +1).
	%\end{equation*}  
	we use \cref{lem:Zeros_F_deg} and \eqref{eq:deg_F} to obtain
	\begin{equation}\label{eq:S_upperbound}
	|\cS| \leq 216q^n \deg \F \leq 1296(L+1)q^n \deg \f.
	\end{equation}
	Combining equations \eqref{eq:S_lowerbound} and \eqref{eq:S_upperbound} gives us
	\begin{equation}\label{eq:estimate_L}
	L \geq \frac{N_{k-r} - 3|\Theta(\Fqp)| - 1296 q^n \deg \f}{1296 q^n \deg \f + 2|\Theta(\Fqp)|}.
	\end{equation}
	By \eqref{eq:number_of_points}, we can estimate the size of $ |\Theta(\Fqn)| $.
	Substituting the lower bound on $ N_{k-r} $ as given in \eqref{eq:bound_Nkr} into \eqref{eq:estimate_L}, we obtain 
	\begin{equation*}
	L(w_{\mb}) \gg   \frac{\min\{ q^{3k/2},\ell/q^8 \}}{ q^n \deg \f } .
	\end{equation*} 
	
\end{proof}	

\section*{Acknowledgements}
	The author was supported by the Austrian Science Fund FWF Project P31762. I am very grateful to L{\'a}szl{\'o} M{\'e}rai for valuable discussions and comments.

\appendix
\section{}
\begingroup
\allowdisplaybreaks
Let $C$ be the hyperelliptic curve defined by \eqref{def:hec} with
$$
f(X) = X^5 + b_1X^4 + b_2X^3 + b_3X^2 + b_4X + b_5 \in \Fq[X],
$$
for the finite field $\Fq$ with characteristic $p \geq 3 $.
\subsection{Defining equations of the Jacobian}
\label{sec:def_eq_J}
Let  
$$
S= \Fq[\X_0, \X_{11},\X_{12},\X_{22},\X_{111},\X_{112},\X_{122},\X_{222},\X] $$ 
be a polynomial ring over field $\Fq$, with characteristic $p \geq 3 $. Following \cite{grant1990formal}, in particular Theorem 2.5, Theorem 2.11 and Corollary 2.15, we define $f_i $ as follows: 
\begin{align*}
f_0 = &\X^2 + \X_{11}^2\X_{12} + b_1\X_{11}^2\X_{22} +   b_2\X_{11}^2\X_{12}\X_{22} - b_3\X_{11}\X_{22}^2 + b_4\X_{12}\X_{22}^2 \\
&-b_5\X_{22}^3 + 2b_1\X\X_{11} -2b_2\X\X_{12} + 2b_3\X\X_{22}+ (b_3-b_1b_2)\X_{11}\X_{12} \\
&+(b_2^2-b_1b_3)\X_{11}\X_{22} + (b_1b_4-b_2b_3-b_5)\X_{12}\X_{22}-b_1b_5\X_{22}^2 \\
& +2(b_1b_3 -b_2^2)\X + (b_1b_4-b_5)\X_{11} +b_2(b_2^2 - b_1b_3 )\X_{12} \\
&(b_3b_4 - b_2b_5)\X_{22} + b_1b_3b_4 - b_2^2b_4 - b_3b_5, \\
f_1 =&2\X - \X_{11}\X_{22} + \X_{12}^2 - b_2\X_{12} + b_4, \\
f_2 =&\X_{112} -\X_{222}\X_{12} + \X_{122}\X_{22}, \\
f_3 =&\X_{111} + \X_{222}\X_{11} + \X_{122}\X_{12} - 2\X_{112}\X_{22} - 2b_1\X_{112} + b_2\X_{122}, \\
f_4 =&\X_{122}^2 - \X_{11}\X_{22}^2 + 2\X\X_{22} + \X_{11}\X_{12} - b_1\X_{11}\X_{22} - b_2\X_{12}\X_{22} \\
& +2b_1\X - b_1b_2\X_{12} + b_4\X_{22} + b_1b_4 - b_5 , \\
f_5 =& \X_{222}^2 - \X_{22}^3 - \X_{12}\X_{22} - b_1\X_{22}^2 - \X_{11} - b_2\X_{22} - b_3 , \\
f_6 =&\X_{122}\X_{222} - \X_{12}\X_{22}^2 + \X -b_2\X_{12} - b_1\X_{12}\X_{22}, \\
f_7 =&\X_{111}^2 - \X_{11}^3 - b_3\X_{11}^2 -b_4\X_{11}\X_{12} + 3b_5\X_{11}\X_{22} + 2b_5\X \\
&+(4b_1b_5 - b_2b_4)\X_{11} - 3b_2b_5\X_{12} + (4b_3b_5-b_4^2)\X_{22} \\
&4b_1b_3b_5 + b_4b_5 - b_1b_4^2 - b_2^2b_5, \\
f_8 =& -\X_{111}\X_{112} + b_1\X_{111}\X_{122} -b_2\X_{112}\X_{122} + b_3\X_{112}\X_{222} \\
&-b_4\X_{122}\X_{222} + b_5\X_{222}^2 -\X^2 -b_1 \X \X_{11} + b_2 \X \X_{12} - b_3 \X \X_{22} \\
&-b_3\X_{11}\X_{12} + b_1b_3\X_{11}\X_{22} - (b_5 + b_1b_4)\X_{12}\X_{22} + 2b_1b_5\X_{22}^2 \\
& -2(b_1b_3 + b_4)\X  + (2b_2b_4 + b_1b_2b_3 +b_1b_5 - b_3^2 -b_1^2b_4)\X_{12} \\
&- 2b_5\X_{11} + 2b_5(b_1^2 - b_2)\X_{22} + b_1b_2b_5 - b_1b_3b_4 - 2b_3b_5 , \\
f_9 =& \X_{122}^2 - \X_{111}\X_{122} + \X_{11}\X - b_3\X_{11}\X_{22} + 2b_4\X_{12}\X_{22} -3b_5\X_{22}^2 \\
&+2b_3\X + (b_1b_4 - b_2b_3 - b_5)\X_{12} - 2b_1 b_5 \X_{22} + b_3b_4 - b_2b_5, \\
f_{10} =&\X_{111}\X_{222} - \X_{112}\X_{122} -2\X\X_{12} + \X_{11}^2 - 2b_1 \X_{11}\X_{12} \\
&+3b_2 \X_{11}\X_{22} - 2b_3 \X_{12}\X_{22} + b_4\X_{22}^2 -5b_2\X +b_3\X_{11} \\
&+ (3b_2^2 - 2b_1b_3)\X_{12} + (b_1b_4 - b_5)\X_{22} - 2b_2b_4, \\
f_{11}=&\X_{122}^2 - \X_{112}\X_{222} + \X_{22}\X + 2\X_{11}\X_{12} - b_1\X_{11}\X_{22} + 2b_1\X \\
&+(b_3-b_1b_2)\X_{12} + b_1b_4 - b_5, \\
f_{12} =&\X_{111}\X_{12} -\X_{112}\X_{11} - b_4\X_{122} + 2b_5\X_{222}, \\
f_{13} =& 2\X_{122}\X_{11} - \X_{112}\X_{12} -\X_{111}\X_{22} - b_2\X_{112} + 2b_3\X_{122} - b_4\X_{222}.
\end{align*}
One can show that $f_0\in \langle f_4,f_5,f_6\rangle$
and the vanishing locus of these polynomials homogenized with respect to the variable $\X_0$ forms a set of defining equations for the Jacobian $J_C$, i.e 
$$
J_C = V(f_1^h,\dots,f_{13}^h) = \{ z \in \mathbb{P}^8(\Fqb) : f_i^h(z) = 0, 1 \leq i \leq 13 \}
$$

\subsection{Addition formulas}
\label{sec:addition_formulas}
For $D = (x_1,y_1) + (x_2,y_2) - 2\cO \in J_C(\Fq)\setminus \Theta(\Fq)$, \eqref{eq:iota} gives us $ \iota(D) $, where the components $z_{jk}, z_{jkl}$ of $\iota(D)$ can be expressed as rational functions in the coordinates $(x_1,y_1)$ and $(x_2,y_2)$. For the sake of completeness, we collect the addition formulas as given in \cite[Theorem 3.3]{grant1990formal} and as explicitly computed in \cite[Appendix A.3]{anupindiLinearComplexitySequences2021}.

\begin{align*}
z_{ij}(Q+R) =& -z_{ij}(Q)-z_{ij}(R) + \frac{1}{4}\left( \frac{q_{i}(Q,R)}{q(Q,R)} \right) \left( \frac{q_{j}(Q,R)}{q(Q,R)} \right) \\
&-\frac{1}{4} \left( \frac{q_{ij}(Q,R)}{q(Q,R)} \right)\\
z_{111}(Q+R)= &-\frac{1}{2}z_{111}(Q)-\frac{1}{2}z_{111}(R) + \frac{3}{16}\frac{q_{1}(Q,R)q_{11}(Q,R)}{q(Q,R)^2}-\frac{1}{16}\frac{q_{111}(Q,R)}{q(Q,R)} \\
&- \frac{1}{8} \left( \frac{q_{1}(Q,R)}{q(Q,R)} \right) ^3 
+ \frac{3}{4}(z_{11}(Q)+z_{11}(R))\frac{q_{1}(Q,R)}{q(Q,R)}, \\
z_{112}(Q+R)= &-\frac{1}{2}z_{112}(Q)-\frac{1}{2}z_{112}(R) + \frac{1}{16}\frac{q_{2}(Q,R)q_{11}(Q,R)}{q(Q,R)^2}\\ &+\frac{1}{8}\frac{q_{1}(Q,R)q_{12}(Q,R)}{q(Q,R)^2} -\frac{1}{16}\frac{q_{112}(Q,R)}{q(Q,R)}
- \frac{1}{8}\frac{q_{2}(Q,R)(q_{1}(Q,R))^2}{q(Q,R)^3} \\  
&+ \frac{3}{8}(z_{11}(Q)+z_{11}(R))\frac{q_{2}(Q,R)}{q(Q,R)} + \frac{3}{8}(z_{12}(Q)+z_{12}(R))\frac{q_{1}(Q,R)}{q(Q,R)} \\
z_{122}(Q+R)= &-\frac{1}{2}z_{122}(Q)-\frac{1}{2}z_{122}(R) + \frac{1}{16}\frac{q_{1}(Q,R)q_{22}(Q,R)}{q(Q,R)^2}\\ &+\frac{1}{8}\frac{q_{2}(Q,R)q_{12}(Q,R)}{q(Q,R)^2} -\frac{1}{16}\frac{q_{122}(Q,R)}{q(Q,R)} \\
&- \frac{1}{8}\frac{q_{1}(Q,R)(q_{2}(Q,R))^2}{q(Q,R)^3}  
+ \frac{3}{4}(z_{12}(Q)+z_{12}(R))\frac{q_{2}(Q,R)}{q(Q,R)} \\
z_{222}(Q+R)= &-\frac{1}{2}z_{222}(Q)-\frac{1}{2}z_{222}(R) + \frac{3}{16}\frac{q_{2}(Q,R)q_{22}(Q,R)}{q(Q,R)^2}-\frac{1}{16}\frac{q_{222}(Q,R)}{q(Q,R)} \\
&- \frac{1}{8} \left( \frac{q_{2}(Q,R)}{q(Q,R)} \right) ^3 
+ \frac{3}{4}(z_{22}(Q)+z_{22}(R))\frac{q_{2}(Q,R)}{q(Q,R)} \\
z(Q+R) =& \frac{1}{2}(z_{11}(Q+R)z_{22}(Q+R) - z_{11}^2(Q+R) + b_2z_{12}(Q+R) - b_4 )
\end{align*}

To evaluate the addition formulas above, we need the following rational functions:
\newcommand{\p}[2]{z_{#1#2}}
\newcommand{\pp}[3]{2z_{#1#2#3}}
\newcommand{\po}{(2z - b_2 z_{12} + b_4)}

\begin{align*}
q(Q,R)= &z_{11}(Q)-z_{11}(R)+z_{12}(Q)z_{22}(R)-z_{12}(R)z_{22}(Q), \\
q_1(Q,R) = &\pp111(Q) - \pp111(R) + \pp112(Q)\p22(R) - \pp112(R)\p22(Q) \\
& +\pp122(R)\p12(Q) - \pp122(Q)\p12(R), \\ 
q_2(Q,R) =& \pp112(Q) - \pp112(R) + \pp122(Q)\p22(R) - \pp122(R)\p22(Q) \\
& +\pp222(R)\p12(Q) - \pp222(Q)\p12(R), \\
q_{11}(Q,R)=& 4b_3 q(Q,R) + 4b_4(\p12(Q)-\p12(R)) + 4(\po(Q)\p12(R))\\
&-4(\po(R)\p12(Q))-8b_5(\p22(Q) - \p22(R)) \\ &+2(\pp112(Q)\pp122(R) - \pp112(R)\pp122(Q)),\\
q_{12}(Q,R) =& 4b_3(\p12(Q)-\p12(R)) + 2b_2(\p12(Q)\p22(R)) \\ &-2b_2(\p12(R)\p22(Q)) - 4(\p11(Q)\p12(R)-\p11(R)\p12(Q))\\
& +2(\po(Q)\p22(R)-\po(R)\p22(Q)) \\
& - 2b_4(\p22(Q) - \p22(R)) + \pp222(R)\pp112(Q)-\pp222(Q)\pp112(R), \\
q_{22}(Q,R) =& 8b_1(\p12(Q)\p22(R)-\p12(R)\p22(Q)) + 4b_2\p12(Q)\\
& -4b_2\p12(R) -8(\p11(Q)\p22(R)-\p11(R)\p22(Q)) \\
& - 4(\po(Q)-\po(R)) \\
& + 2(\pp122(Q)\pp222(R) -\pp122(R)\pp222(Q)),\\
q_{111}(Q,R) =& 4b_3q_1(Q,R) \\
& + 4(\pp111(Q)\p22(Q)\p12(R)-\pp111(R)\p22(R)\p12(Q)) \\
& + \pp122(R)(2\p12(Q)(6\p11(Q) - 2\p11(R)+4b_3)-4b_4\p22(Q))\\
& - \pp122(Q)(2\p12(R)(6\p11(R) - 2\p11(Q)+4b_3)-4b_4\p22(R))\\
& + \pp112(Q)(\p12(R)(12\p12(R) - 8\p12(Q)+4b_2)+4b_4)\\
& - \pp112(R)(\p12(Q)(12\p12(Q) - 8\p12(R)+4b_2)+4b_4)\\
q_{112}(Q,R) = &~\pp222(Q)
\left(4 \p11(Q) \p12(R) - 4 \p12(R) b_{3} - 8 b_{5}
\right) \\ 
&+\pp112(Q)
\left(- 4 \p11(R) + 4 \p12(R) \p22(Q) + \p12(R) \left(12 \p22(R) + 8 b_{1}\right)
\right) \\ 
&+\pp112(R)
\left(4 \p11(Q) + \p12(Q) \left(- 12 \p22(Q) - 4 \p22(R) - 8 b_{1}\right) - 4 b_{3}
\right) \\ 
&+\pp122(Q)
(- 8 \p11(R) \p22(R) - 8 \p12(Q) \p12(R) - 4 \p12(R)^{2}  \\
&+ 4 \p22(R) b_{3} + 4 b_{4} - 4 \p12(R) b_{2} ) \\ 
&+\pp122(R)
\left(8 \p11(Q) \p22(Q) + 4 \p12(Q)^{2} + \p12(Q) \left(8 \p12(R) + 4 b_{2}\right) \right.\\
&- \left. 4 \p22(Q) b_{3} - 4 b_{4} \right) + \pp112(Q)4 b_{3}\\ 
&+\pp222(R)
\left(\p12(Q) \left(- 4 \p11(R) + 4 b_{3}\right) + 8 b_{5}
\right) \\
q_{122}(Q,R) = &~\pp112(R)
\left(- 6 \p22(Q)^{2} + \p22(Q) \left(- 2 \p22(R) - 4 b_{1}\right) - 2 b_{2}
\right) \\ 
&+\pp122(R)
\left(- 4 \p11(Q) + \p22(Q) \left(4 \p12(R) - 2 b_{2}\right) - 4 b_{3}
\right) \\ 
&+\pp222(Q)
\left(2 \p11(Q) \p22(R) - 4 \p11(R) \p22(R) - 2 \p12(R)^{2}   
\right) \\ 
&+\pp112(Q)
\left(2 \p22(Q) \p22(R) + 6 \p22(R)^{2} + 4 \p22(R) b_{1} + 2 b_{2}
\right) \\ 
&+\pp222(R)
\left(4 \p11(Q) \p22(Q) - 2 \p11(R) \p22(Q) + 2 \p12(Q)^{2} 
\right) \\ 
&+\pp122(Q)
\left(4 \p11(R) - 4 \p12(Q) \p22(R) + 2 \p22(R) b_{2} + 4 b_{3}
\right) \\
&-\pp222(Q) (2 b_{4} + 4 \p12(R) b_{2}) \\
&+ \pp222(R) (2 b_{4} + 4 \p12(Q) b_{2}) \\
q_{222}(Q,R) = &~\pp222(R)
\left(- 12 \p11(Q) + 4 \p11(R) + \p12(Q) \left(12 \p22(Q) + 16 b_{1}\right)
\right) \\ 
&+\pp122(R)
\left(- 8 \p12(Q) - 8 \p12(R) - 12 \p22(Q)^{2} - 16 \p22(Q) b_{1} - 8 b_{2}
\right) \\ 
&+\pp112(Q)
\left(- 4 \p22(Q) - 8 \p22(R)
\right) \\ 
&+\pp222(Q)
\left(- 4 \p11(Q) + 12 \p11(R) + \p12(R) \left(- 12 \p22(R) - 16 b_{1}\right)
\right) \\ 
&+\pp112(R)
\left(8 \p22(Q) + 4 \p22(R)
\right) \\ 
&+\pp122(Q)
\left(8 \p12(Q) + 8 \p12(R) + 12 \p22(R)^{2} + 16 \p22(R) b_{1} + 8 b_{2}
\right)
\end{align*}
\endgroup

\bibliographystyle{amsplain}
\bibliography{Bibliography}
\end{document}